OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x^2 + ___x + 49

OpenStudy (campbell_st):

the perfect square is (2x + 7) the middle term is 2*2x*7 = 28x so the blank should be 28

OpenStudy (diyadiya):

@careless850 Do you remember how to find Discriminant?

OpenStudy (diyadiya):

\( Discriminant = b^2 - 4ac \) Since its a perfect square trinmial Discriminant will be 0 (zero) \[b^2 - 4ac=0\]\[b^2 - 4 \times 4 \times 49 =0\]\[b^2-784=0\]Add 784 on both sides \[b^2-784 + 784 = 784\]\[b^2 = 784\]Now you take square root on both sides\[\sqrt{b^2}-\sqrt{784}\]\[b=\sqrt{784}\]\[b=28\]

OpenStudy (anonymous):

... wow I feel so dumb.